OpenStudy (anonymous):

find the distance between the pair of points (-12, -4) and (7, 15)

OpenStudy (anonymous):

use distance formula i.e \[\sqrt{(x1-x0)^{2}+(y1-y0)^{2}}\] where x1=7,x0=-12 and y0=-4,y1=15

OpenStudy (anonymous):

SO ITLL LOOK LIKE (71-12)^2+(-41-150)^2?

OpenStudy (anonymous):

(71-(-12))^2+(-41-150)^2

OpenStudy (anonymous):

okay...so I just solve that and that's it?

hartnn (hartnn):

its not 71!!

OpenStudy (anonymous):

nice to hear that you have solve it :)

hartnn (hartnn):

(7-(-12))^2 + (15-(-4))^2 then take square root.

OpenStudy (anonymous):

yes its a typing mistak thanks for correction

OpenStudy (anonymous):

okay.so it's not 71? mominanisar probablly made a mistake and didn't see that. It's okay :) thanks for helping me yal!!

OpenStudy (anonymous):

I got 722...is that right? lol...I mean the answer before I take the square root of it

hartnn (hartnn):

yes.

OpenStudy (anonymous):

okay thank you!! I get it now!!

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]